### v1

@@ 2.1 Exact and stable recovery
Proof 2.12.
The proof is similar to the one of [yin2015minimization, Theorem 2.2] and so we omit it here.
2.2 Sparsity and Extremality under NSP
This subsection primarily investigates the relationship between sparsity and extremality of CDCEN and DCEN under the null space property (NSP) condition.
Definition 2.13.
For any subset $\Lambda\subset\{1,2,\cdots,n\}$ with $|\Lambda|=s$, a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ is said to satisfy the DCEN stable null space property (DCEN-NSP) of order $s$ if $\|\bm{h}_{\Lambda}\|_{1}\leq\kappa\|\bm{h}_{\Lambda^{c}}\|_{1}$ for all $\bm{h}\in\mathrm{ker}(\bm{A})$, where $\kappa\in(0,1)$ is a constant.
Theorem 2.14.
Let $\bm{x}^{*}\in\mathbb{R}^{n}$ be a feasible point of the problem (1) with $r:=\|\bm{x}^{*}\|_{2}$. Denote the support set of $\bm{x}^{*}$ by $\Lambda_{*}:=\mathrm{supp}(\bm{x}^{*})=\{i\mid x_{i}^{*}\neq 0\}$. For any nonzero vector $\bm{h}\in\ker(\bm{A})\setminus\{\bm{0}\}$ with $\bm{h}=\bm{h}_{\Lambda_{*}}+\bm{h}_{\Lambda_{*}^{c}}$, where $\Lambda_{*}^{c}$ is the complement of $\Lambda_{*}$. Let $s_{1}:=\|\bm{h}_{\Lambda_{*}}\|_{1}$, $s_{2}:=\|\bm{h}_{\Lambda_{*}^{c}}\|_{1}$ and $\kappa(r):=\gamma(1-\alpha)/[\gamma(1+\alpha)+2(1-\gamma)r]\in(0,1)$. If there exists a constant $\kappa\in(0,\kappa(r))$ such that the DCEN-NSP condition $s_{1}<\kappa s_{2}$ holds for all $\bm{h}\in\ker(\bm{A})\setminus\{\bm{0}\}$, then $\bm{x}^{*}$ is the unique minimizer of the problem (1).
Proof 2.15.
To show that $\bm{x}^{*}$ is the unique minimizer of (1), it suffices to prove that $r_{\gamma}(\bm{x}^{*}+\tau\bm{h})>r_{\gamma}(\bm{x}^{*})$ holds for any $\bm{h}\in\ker(\bm{A})$ with $\bm{h}\neq\bm{0}$ and $\tau>0$. For any $\bm{h}\in\ker(\bm{A})\setminus\{\bm{0}\}$ and any $\tau>0$, consider $\bm{x}=\bm{x}^{*}+\tau\bm{h}$. Since $\bm{Ax}^{*}=\bm{b}$ and $\bm{Ah}=\bm{0}$, we have $\bm{A}(\bm{x}^{*}+\tau\bm{h})=\bm{A}\bm{x}^{*}+\tau\bm{A}\bm{h}=\bm{b}$, which means that $\bm{x}$ is still feasible. Let the increment of the objective function as $\Delta(\tau):=r_{\gamma}(\bm{x}^{*}+\tau\bm{h})-r_{\gamma}(\bm{x}^{*})$. Then
(20) $\Delta(\tau)$ $=\gamma\left(\|\bm{x}^{*}+\tau\bm{h}\|_{1}-\|\bm{x}^{*}\|_{1}\right)-\gamma\alpha\left(\|\bm{x}^{*}+\tau\bm{h}\|_{2}-\|\bm{x}^{*}\|_{2}\right)$
$+(1-\gamma)\left(\|\bm{x}^{*}+\tau\bm{h}\|_{2}^{2}-\|\bm{x}^{*}\|_{2}^{2}\right).$
Since $\bm{x}_{\Lambda_{*}^{c}}^{*}=0$, it is easy to have $\|\bm{x}^{*}+\tau\bm{h}\|_{1}=\|\bm{x}_{\Lambda_{*}}^{*}+\tau\bm{h}_{\Lambda_{*}}\|_{1}+\tau\|\bm{h}_{\Lambda_{*}^{c}}\|_{1}$. By the fact $\|\bm{x}_{\Lambda_{*}}^{*}+\tau\bm{h}_{\Lambda_{*}}\|_{1}\geq\|\bm{x}_{\Lambda_{*}}^{*}\|_{1}-\tau\|\bm{h}_{\Lambda_{*}}\|_{1}=\|\bm{x}^{*}\|_{1}-\tau\|\bm{h}_{\Lambda_{*}}\|_{1}$, it follows that $\|\bm{x}^{*}+\tau\bm{h}\|_{1}-\|\bm{x}^{*}\|_{1}\geq\tau\|\bm{h}_{\Lambda_{*}^{c}}\|_{1}-\tau\|\bm{h}_{\Lambda_{*}}\|_{1}=\tau(s_{2}-s_{1})$ and so
(21) $\gamma\left(\|\bm{x}^{*}+\tau\bm{h}\|_{1}-\|\bm{x}^{*}\|_{1}\right)\geq\gamma\tau(s_{2}-s_{1}).$
For the term $\gamma\alpha(\|\bm{x}^{*}+\tau\bm{h}\|_{2}-\|\bm{x}^{*}\|_{2})$, by the triangle inequality, we have $\|\bm{x}^{*}+\tau\bm{h}\|_{2}-\|\bm{x}^{*}\|_{2}\leq\tau\|\bm{h}\|_{2}=\tau\sigma$ and so
(22) $\gamma\alpha(\|\bm{x}^{*}+\tau\bm{h}\|_{2}-\|\bm{x}^{*}\|_{2})\geq-\gamma\alpha\tau\sigma,$
where $\sigma:=\|\bm{h}\|_{2}$. For the quadratic term $(1-\gamma)\|\bm{x}\|_{2}^{2}$, its expansion gives $\|\bm{x}^{*}+\tau\bm{h}\|_{2}^{2}-\|\bm{x}^{*}\|_{2}^{2}=2\tau\bm{x}^{*\top}\bm{h}+\tau^{2}\|\bm{h}\|_{2}^{2}$. It follows that $(1-\gamma)(\|\bm{x}^{*}+\tau\bm{h}\|_{2}^{2}-\|\bm{x}^{*}\|_{2}^{2})=2(1-\gamma)\tau\bm{x}^{*\top}\bm{h}+(1-\gamma)\tau^{2}\sigma^{2}$. Noting that $\bm{x}^{*\top}\bm{h}=\bm{x}_{\Lambda_{*}}^{*\top}\bm{h}_{\Lambda_{*}}$, by the Cauchy-Schwarz inequality we have $|\bm{x}_{\Lambda_{*}}^{*\top}\bm{h}_{\Lambda_{*}}|\leq\|\bm{x}_{\Lambda_{*}}^{*}\|_{2}\|\bm{h}_{\Lambda_{*}}\|_{2}\leq r\|\bm{h}_{\Lambda_{*}}\|_{1}=rs_{1}$, and using $\|\bm{h}_{\Lambda_{*}}\|_{2}\leq\|\bm{h}_{\Lambda_{*}}\|_{1}$, we further obtain
(23) $(1-\gamma)(\|\bm{x}^{*}+\tau\bm{h}\|_{2}^{2}-\|\bm{x}^{*}\|_{2}^{2})\geq-2(1-\gamma)\tau rs_{1}+(1-\gamma)\tau^{2}\sigma^{2}.$
Combining the inequalities (21), (22), and (23), we have $\Delta(\tau)=\geq\gamma\tau(s_{2}-s_{1})-\gamma\alpha\tau\sigma-2(1-\gamma)\tau rs_{1}+(1-\gamma)\tau^{2}\sigma^{2}$. This leads to a quadratic lower bound $\Delta(\tau)\geq\tau\Big(\gamma(s_{2}-s_{1})-\gamma\alpha\sigma-2(1-\gamma)rs_{1}\Big)+(1-\gamma)\tau^{2}\sigma^{2}$. Since $\sigma=\|\bm{h}\|_{2}\leq\|\bm{h}\|_{1}=\|\bm{h}_{\Lambda_{*}}\|_{1}+\|\bm{h}_{\Lambda_{*}^{c}}\|_{1}=s_{1}+s_{2}$, we have
(24) $\gamma(s_{2}-s_{1})-\gamma\alpha\sigma-2(1-\gamma)rs_{1}\geq\gamma(s_{2}-s_{1})-\gamma\alpha(s_{1}+s_{2})-2(1-\gamma)rs_{1}.$
Substituting $s_{1}<\kappa s_{2}$ into the inequality (24) gives $\gamma(1-\alpha)s_{2}-(\gamma(1+\alpha)+2(1-\gamma)r)s_{1}\geq\gamma(1-\alpha)s_{2}-(\gamma(1+\alpha)+2(1-\gamma)r)\kappa(r)s_{2}=0$. Thus, the right-hand side of (20) is strictly positive, implying $\Delta(\tau)>0$ for all $\tau>0$ and $\bm{h}\in\ker(\bm{A})\setminus\{\bm{0}\}$.
For any other feasible point $\bm{x}^{\prime}\neq\bm{x}^{*}$, let $\bm{h}=\bm{x}^{\prime}-\bm{x}^{*}\in\ker(\bm{A})$ and choose $\tau=1$ in (20). Then $r_{\gamma}(\bm{x}^{\prime})-r_{\gamma}(\bm{x}^{*})=\Delta(1)>0$ and so $\bm{x}^{*}$ is the unique global minimizer of (1).
Corollary 2.16.
Let $\bm{x}^{*}$ be a strict local minimizer of problem (1) with $r=\|\bm{x}^{*}\|_{2}$, and denote its support by $\Lambda_{*}=\mathrm{supp}(\bm{x}^{*})$. If the matrix $\bm{A}$ satisfies
(25) $\|\bm{h}_{\Lambda_{*}}\|_{1}<\kappa\,\|\bm{h}_{\Lambda_{*}^{c}}\|_{1},\quad\kappa\in\left(0,\kappa(r)\right),$
for all nonzero $\bm{h}\in\ker(\bm{A})$, then the submatrix $\bm{A}_{\Lambda_{*}}$ formed by the columns indexed by $\Lambda_{*}$ is of full column rank, i.e., $\mathrm{rank}(\bm{A}_{\Lambda_{*}})=|\Lambda_{*}|.$
Proof 2.17.
We prove the corollary by contradiction. If the submatrix $\bm{A}_{\Lambda_{*}}$ is not of full column rank, there exists a nonzero vector $\bm{u}\in\mathbb{R}^{|\Lambda_{*}|}$ such that $\bm{A}_{\Lambda_{*}}\bm{u}=0$. Extend $\bm{u}$ to a vector $\bm{h}\in\mathbb{R}^{n}$ by setting $\bm{h}_{\Lambda_{*}}=\bm{u}$ and $\bm{h}_{\Lambda_{*}^{c}}=\bm{0}$. For this $\bm{h}$, we have $\|\bm{h}_{\Lambda_{*}}\|_{1}>0$ and $\|\bm{h}_{\Lambda_{*}^{c}}\|_{1}=0$, which contradicts. Hence, $\bm{A}_{\Lambda_{*}}$ must have linearly independent columns.
Remark 2.18.
This corollary indicates that the condition (25) not only provides a sufficient criterion for strict local optimality but also ensures algebraic consistency between the measurement matrix $\bm{A}$ and the sparsity pattern of the solution: the columns corresponding to the support must be linearly independent; otherwise, a nonzero vector existing only on the support would lie in $\ker(\bm{A})$, violating the inequality condition.
Corollary 2.19.
Let $\bm{x}^{*}$ be a strict local minimizer of problem (2). If the matrix $\bm{A}$ satisfies (25) for all nonzero $\bm{h}\in\ker(\bm{A})$, then the submatrix $\bm{A}_{\Lambda_{*}}$ is of full column rank.
Proof 2.20.
We claim that $\bm{x}^{*}$ is also a local minimizer of the following constrained problem
$\min_{\bm{x}\in\mathbb{R}^{n}}\gamma(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2})+(1-\gamma)\|\bm{x}\|_{2}^{2}\quad\text{subject to}\quad\bm{Ax}=\bm{Ax}^{*}=\bm{b}.$
Suppose not. Then $\forall\tilde{r}>0$, there exists $\bm{\tilde{x}}\in\mathcal{B}_{\tilde{r}}(\bm{x}^{*})$ such that $\bm{A\tilde{x}}=\bm{Ax}^{*}$ and $\gamma(\|\bm{\tilde{x}}\|_{1}-\alpha\|\bm{\tilde{x}}\|_{2})+(1-\gamma)\|\bm{\tilde{x}}\|_{2}^{2}<\gamma(\|\bm{x}^{*}\|_{1}-\alpha\|\bm{x}^{*}\|_{2})+(1-\gamma)\|\bm{x}^{*}\|_{2}^{2}$. This implies $\frac{1}{2}\|\bm{A\tilde{x}}-\bm{b}\|_{2}^{2}+\lambda(\gamma(\|\bm{\tilde{x}}\|_{1}-\alpha\|\bm{\tilde{x}}\|_{2})+(1-\gamma)\|\bm{\tilde{x}}\|_{2}^{2})<\frac{1}{2}\|\bm{Ax}^{*}-\bm{b}\|_{2}^{2}+\lambda(\gamma(\|\bm{x}^{*}\|_{1}-\alpha\|\bm{x}^{*}\|_{2})+(1-\gamma)\|\bm{x}^{*}\|_{2}^{2})$. Thus for any $\tilde{r}>0$, we actually find a $\bm{\tilde{x}}\in\mathcal{B}_{\tilde{r}}(\bm{x}^{*})$ yielding a smaller objective of (2) than $\bm{x}^{*}$, which leads to a contradiction because $\bm{x}^{*}$ is assumed to be a local minimizer. Thus the claim is validated. Using the claim above and Corollary 2.16, we have that the columns of $\bm{A}_{\Lambda_{*}}$ are linearly independent.
3 Proximal Operator

### v2

@@ II-A Exact and stable recovery
Proof.
The proof is similar to the one of [41, Theorem 2.2] and so we omit it here. ∎
III Consistency Theory for DCEN
In this section, we study the consistency theory of DCEN. Let $\bar{\bm{x}}$ be the ground true $s$-sparse solution with support $\mathcal{S}:=\{i:\bar{x}_{i}\neq 0\}$ and $s:=|\mathcal{S}|$. We define the level set of the objective function at $\bar{\bm{x}}$ as $\text{lev}_{h}(\bar{\bm{x}}):=\{\bm{x}\in\mathbb{R}^{n}:h(\bm{x})\leq h(\bar{\bm{x}})\}$. Associated with the proposed regularization, we introduce an extended notion of the Restricted Eigenvalue Condition (REC).
[…]
$\begin{split}\phi_{\beta}(s,t):=\min\bigg\{&\frac{\|\bm{A}\bm{x}\|_{2}}{\|\bm{x}_{\mathcal{J}}\|_{2}}:|\mathcal{I}|\leq s,\|\bm{x}_{\mathcal{I}^{c}}\|_{1}\leq\|\bm{x}_{\mathcal{I}}\|_{1}\\ &\quad+\beta\|\bm{x}\|_{2},\mathcal{J}=\mathcal{I}\cup\mathcal{I}(\bm{x};t)\bigg\}>0.\end{split}$ (20)
where $\mathcal{I}(\bm{x};t)$ denotes the index set of the first $t$ largest coordinates in absolute value of $\bm{x}$ in $\mathcal{I}^{c}$.
The following theorem provides the oracle inequality and recovery bound for DCEN.
Theorem 9.
[…]
  1. (i)
    The oracle inequality holds:
    $\ell(\bm{x}^{*})+\lambda\gamma\bigl(\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\bigr)\leq\frac{2\Gamma^{2}}{\phi_{\beta}^{2}(s,s)}.$ (21)
  2. (ii)
    The recovery bound holds:
    $\|\bm{x}^{*}-\bar{\bm{x}}\|_{2}^{2}\leq\frac{\Gamma^{2}}{\phi_{\beta}^{4}(s,s)}\left(4+\frac{\Gamma^{2}}{4s\lambda^{2}\gamma^{2}(1-\alpha)^{2}}\right).$ (22)
Proof.
(i) Let $\bm{x}^{*}\in\text{lev}_{h}(\bar{\bm{x}})$. Then $h(\bm{x}^{*})\leq h(\bar{\bm{x}})$. Denoting the residual as $\hat{\bm{x}}:=\bm{x}^{*}-\bar{\bm{x}}$ and recalling $\bm{b}=\bm{A}\bar{\bm{x}}$, we have
$\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\lambda r(\bm{x}^{*})\leq\lambda r(\bar{\bm{x}}).$ (23)
By the subadditivity of norms and the fact that $\bar{\bm{x}}_{\mathcal{S}^{c}}=\bm{0}$, which implies $\bm{x}^{*}_{\mathcal{S}^{c}}=\hat{\bm{x}}_{\mathcal{S}^{c}}$.Then, we have
$\|\bar{\bm{x}}\|_{1}-\|\bm{x}^{*}\|_{1}$ $=\|\bar{\bm{x}}_{\mathcal{S}}\|_{1}-(\|\bm{x}^{*}_{\mathcal{S}}\|_{1}+\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1})$ (24)
$\leq\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}-\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{1}$
and $\alpha\|\bm{x}^{*}\|_{2}\geq-\alpha(\|\bm{x}^{*}_{\mathcal{S}}\|_{2}+\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2})=-\alpha\|\bm{x}^{*}_{\mathcal{S}}\|_{2}-\alpha\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{2}$. Moreover, the reverse triangle inequality yields $\|\bm{x}^{*}_{\mathcal{S}}\|_{2}-\|\bar{\bm{x}}\|_{2}\leq\|\bm{x}^{*}_{\mathcal{S}}-\bar{\bm{x}}_{\mathcal{S}}\|_{2}=\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}$, which leads to $\alpha\|\bm{x}^{*}_{\mathcal{S}}\|_{2}-\alpha\|\bar{\bm{x}}\|_{2}\leq\alpha\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}$. Applying the Cauchy-Schwarz inequality, one has $\|\bar{\bm{x}}\|_{2}^{2}-\|\bm{x}^{*}\|_{2}^{2}=-2\langle\bar{\bm{x}},\hat{\bm{x}}\rangle-\|\hat{\bm{x}}\|_{2}^{2}\leq 2\|\bar{\bm{x}}\|_{2}\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}-\|\hat{\bm{x}}\|_{2}^{2}$. Substituting these relations into (23) and rearranging yields
$\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\lambda(1-\gamma)\|\hat{\bm{x}}\|_{2}^{2}+\lambda\gamma\bigl(\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{2}\bigr)$ (25)
$\quad\leq\lambda\gamma\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+\lambda\gamma\alpha\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}+2\lambda(1-\gamma)\|\bar{\bm{x}}\|_{2}\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}.$
It follows from $\|\bar{\bm{x}}\|_{2}\leq R$ and $\gamma\geq\frac{2R}{2R+\sqrt{s}}$ that
$2\lambda(1-\gamma)\|\bar{\bm{x}}\|_{2}\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}\leq 2\lambda(1-\gamma)R\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}\leq\lambda\gamma\sqrt{s}\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}.$ (26)
Substituting (26) back into the RHS of (25), we obtain
$\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\lambda(1-\gamma)\|\hat{\bm{x}}\|_{2}^{2}+\lambda\gamma\bigl(\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{2}\bigr)$ (27)
$\leq\lambda\gamma\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+\lambda\gamma(\alpha+\sqrt{s})\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}.$
Then, one has $\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{1}\leq\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+\alpha\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{2}+(\alpha+\sqrt{s})\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}$. Since both $\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{2}\leq\|\hat{\bm{x}}\|_{2}$ and $\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}\leq\|\hat{\bm{x}}\|_{2}$, it strictly follows that
$\|\hat{\bm{x}}_{\mathcal{S}^{c}}\|_{1}\leq\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+(2\alpha+\sqrt{s})\|\hat{\bm{x}}\|_{2}=\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+\beta\|\hat{\bm{x}}\|_{2}.$
Thus, $\hat{\bm{x}}$ geometrically falls into the restricted cone, and the condition $\phi_{\beta}(s,s)$ is applicable, implying $\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}\leq\frac{1}{\phi_{\beta}(s,s)}\|\bm{A}\hat{\bm{x}}\|_{2}$. It follows from $\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}\leq\sqrt{s}\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}$ that
$\lambda\gamma(\|\hat{\bm{x}}_{\mathcal{S}}\|_{1}+(\alpha+\sqrt{s})\|\hat{\bm{x}}_{\mathcal{S}}\|_{2})\leq\lambda\gamma(2\sqrt{s}+\alpha)\|\hat{\bm{x}}_{\mathcal{S}}\|_{2}.$
Dropping the $\lambda(1-\gamma)\|\hat{\bm{x}}\|_{2}^{2}$ term on the LHS of (27) gives
$\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\lambda\gamma\bigl(\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\bigr)\leq\frac{\Gamma}{\phi_{\beta}(s,s)}\|\bm{A}\hat{\bm{x}}\|_{2}.$ (28)
Applying the inequality $ab\leq\frac{1}{2}a^{2}+\frac{1}{2}b^{2}$ to the RHS of (28) with $a=\|\bm{A}\hat{\bm{x}}\|_{2}$ and $b=\frac{\Gamma}{\phi_{\beta}(s,s)}$, we obtain
$\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\lambda\gamma\bigl(\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\bigr)\leq\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}+\frac{\Gamma^{2}}{2\phi_{\beta}^{2}(s,s)}.$
Subtracting $\frac{1}{2}\|\bm{A}\hat{\bm{x}}\|_{2}^{2}$ temporarily gives $\lambda\gamma\bigl(\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\bigr)\leq\frac{\Gamma^{2}}{2\phi_{\beta}^{2}(s,s)}$. Re-substituting this bound back into (28) yields the oracle inequality (21).
(ii) Since $\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\leq\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}$, we have the strict lower bound for the nonconvex penalty $\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}-\alpha\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{2}\geq(1-\alpha)\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}$. Combining this with (21), we explicitly bound the tail $\ell_{1}$ norm
$\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}\leq\frac{\Gamma^{2}}{2\lambda\gamma(1-\alpha)\phi_{\beta}^{2}(s,s)}.$
Define the extended support $\mathcal{N}:=\mathcal{S}\cup\mathcal{S}(\bm{x}^{*};s)$. By Lemma 7 in [17], one has
$\|\hat{\bm{x}}_{\mathcal{N}^{c}}\|_{2}\leq\frac{1}{\sqrt{s}}\|\bm{x}^{*}_{\mathcal{S}^{c}}\|_{1}\leq\frac{\Gamma^{2}}{2\sqrt{s}\lambda\gamma(1-\alpha)\phi_{\beta}^{2}(s,s)}.$ (29)
[…]
$\|\hat{\bm{x}}_{\mathcal{N}}\|_{2}\leq\frac{\|\bm{A}\hat{\bm{x}}\|_{2}}{\phi_{\beta}(s,s)}\leq\frac{2\Gamma}{\phi_{\beta}^{2}(s,s)},$ (30)
where the last inequality utilizes the bound $\|\bm{A}\hat{\bm{x}}\|_{2}\leq\frac{2\Gamma}{\phi_{\beta}(s,s)}$ derived from (28). Summing the squared norms (29) and (30) due to disjoint support, we have
$\|\bm{x}^{*}-\bar{\bm{x}}\|_{2}^{2}$ $=\|\hat{\bm{x}}_{\mathcal{N}}\|_{2}^{2}+\|\hat{\bm{x}}_{\mathcal{N}^{c}}\|_{2}^{2}$
$\leq\frac{\Gamma^{2}}{\phi_{\beta}^{4}(s,s)}\left(4+\frac{\Gamma^{2}}{4s\lambda^{2}\gamma^{2}(1-\alpha)^{2}}\right),$
which establishes (22). The proof is complete. ∎
Remark 10.
Compared to the $\ell_{1-2}$ model ($\alpha=1,\gamma=1$), the proposed Theorem 9 mathematically demonstrates a highly sophisticated trade-off:
  1. 1.
    Elimination of Mathematical Singularity: By restricting $\alpha\in(0,1)$ and adding $\ell_{2}^{2}$, DCEN provides a strict coercivity bound. This completely eliminates the fatal mathematical singularity (the $\frac{1}{\sqrt{s}-2}$ blow-up) present in the original $\ell_{1-2}$ theory, guaranteeing a globally continuous and structurally stable recovery bound for any arbitrary sparsity level $s\geq 1$.
  2. 2.
    Enhanced Robustness via the Grouping Effect: The incorporation of the $\ell_{2}^{2}$ penalty (controlled by $1-\gamma$) structurally regularizes the Hessian matrix. This effectively expands the restricted cone (allowing a larger $\beta$), granting the model immense robustness against highly coherent design matrices where pure nonconvex methods fail.
IV Proximal Operator